OpenStudy (anonymous):

If the voltage, V, in an electric circuit is held constant, the current, I, is inversely proportional to the resistance, R. If the current is 360 milliamperes when the resistance is 2 ohms, find the current when the resistance is 12 ohms.

OpenStudy (anonymous):

U = R * i U = 2 * 360 = 720 mV U = R * i 720 = 12 * i i = 60mA